OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (3, 4) and is parallel to the line y = –x + 6?

OpenStudy (aroub):

y=mx+b , m= -1 , (3,4) 4=-1(3)+b 4=-3+b 7=b y=-x+7